OpenStudy (anonymous):

Each term of a certain sequence is greater than the term before it. The difference between any two consecutive terms in the sequence is always the same number. If the fourth and the ninth term of the sequence are 121 and 176 respectively, what is the value of the second term?

OpenStudy (anonymous):

thxs for helping me mebs. ;__; means so much. i really want to improve my SAT score D':

OpenStudy (anonymous):

a1 a2 a3... its an arithmatic sequence so.... recall that \[an = a1 + (n-1)d\] you know that that 121 = a4 + (4-1)d you also know that 176 = a9 +(9-1)d

OpenStudy (anonymous):

and a4 and a9 are the same term...

OpenStudy (anonymous):

so know we can say that : 176 - 8d = 121 - 3d right? you follow?

OpenStudy (anonymous):

wait how are they the same D:

OpenStudy (anonymous):

because the first term is the same...

OpenStudy (anonymous):

remember that the same sequence has the same first term... right.. please tell you studied geometric and arithmetic sequences.... you did right?

OpenStudy (anonymous):

yes, but i forgot it :(

OpenStudy (anonymous):

that's ok lets review.... here is a sequence: 2 , 4 , 6 , 8 , 10 here is another sequence: 2 , 4, 8, 16, 32 what is it and what Is the difference.

OpenStudy (anonymous):

First sequence: arithmetic, d = 2 2: geometric, r = 2

OpenStudy (anonymous):

good good... now for this one we are doing the general formula an = a1 + (n-1)d then a1 is the first term for the arithmetic sequence... so that's why I said a4 and a9 are the same...

OpenStudy (anonymous):

when you said, "you know that that 121 = a4 + (4-1)d you also know that 176 = a9 +(9-1)d" does it work if i do this to it 121 = 121 + (4-1)d 176 = 176 + (9-1)d ?

OpenStudy (anonymous):

no that's not right 121 is the 4th term not the 1st tem... and 176 is the 9th term not the 1st term...

OpenStudy (anonymous):

its 121 = a1 +(3)d 176 = a1 +(8)d

OpenStudy (anonymous):

ohhhhh oh oh ohhhh xD omg, i see it. thx thx :D sorry for my brain. xD

OpenStudy (anonymous):

good nor rearrange for a1 and now you get a1 = 121 -3d a1 = 176 -8d alright?

OpenStudy (anonymous):

yes :)!

OpenStudy (anonymous):

now you can solve this simple equation of 176 - 8d = 121 -3d right? whats the answer?

OpenStudy (anonymous):

d is 11 :D!~~~

OpenStudy (anonymous):

good now find what a is

OpenStudy (anonymous):

then add 11 to it to find the second term and your done!

OpenStudy (anonymous):

BD Thanks again a2 = 99

OpenStudy (anonymous):

well done! yw